a. Purchased raw materials on credit, $240,000. b. Materials requisitions record use of the following materials for the monthRequired: 1. Prepare a job cost sheet for each job worked on during the month. Job No. 136 Job No. 137 Job No. 138 Job No. 132. Prepare journal entries to record the events and transactions a through i. View transaction list Journal entry worksheet <3. Prepare T-accounts for each of the following general ledger accounts, each of which started the month with a zero balance:4. Prepare a report showing the total cost of each job in process and prove that the sum of their costs equals the Work in Pr
Answer #1

1)

Job No 136 Job No 137 Job No 138 Job No 139 Job No 140
Materials $48000 $32500 $19800 $23400 $7000
Labor 12200 10500 37900 39400 3600
Overhead (12200*200%)= 24400 (10500*200%)= 21000 (37900*200%)= 75800 (39400*200%)= 78800 (3600*200%)= 7200
Total cost $84600 $64000 $133500 $141600 $17800

2)

No. Date Account titles and explanation Debit Credit
1. a Raw materials inventory $240000
Account payable $240000
(To record raw materials purchased on account)
2. b Work in process inventory $130700
Factory overhead $19500
Raw materials inventory $150200
(To record raw materials used)
3. c Factory overhead $15750
Cash $15750
(To record cash paid)
4. d Work in process inventory $103600
Factory overhead $26500
Cash $130100
(To record direct and indirect labor paid in cash)
5. e Work in process inventory ($24400+75800+78800) $179000
Factory overhead $179000
(To record factory overhead applied of Jobs 136, 138 and 139)
6. f Finished goods inventory ($84600+133500+141600) $359700
Work in process inventory $359700
(To record transferred of Job 136, 138 and 139 to finished goods)
7. g Account receivable $545000
Sales $545000
(To record sales)
8. Cost of goods sold ($84600+133500) $218100
Finished goods inventory $218100
(To record cost of goods sold)
9. h Factory overhead $156000
Accumulated depreciation- Factory building $70000
Accumulated depreciation- Factory equipment $38500
Prepaid insurance $11000
Property tax payable $36500
(To record actual factory overhead)
10. i Work in process inventory ($10500+3600)*200% $28200
Factory overhead $28200
(To record overhead applied of Job 137 and 140)

3)

Raw materials Inventory Work in process Inventory
a $240000 150200 b b $130700 359700 f
d 103600
e 179000
End. bal. $89800 i 28200
End. bal. $81800
Factory overhead Finished goods Inventory
b $19500 179000 e f $359700 218100 g
c 15750 28200 i
d 26500
h 156000 End. bal. $141600
End. bal. $10550
Cost of Goods Sold
g $218100
End. bal. $218100

4)

Report of Job Costs
Work in process Inventory
Job 137 $64000
Job 140 17800
Balance $81800
Finished goods Inventory
Job 139 $141600
Balance $141600
Cost of Goods Sold
Job 136 $84600
Job 138 133500
Balance $218100